Henry drove 800 miles in X hours on Monday.
A. Write a rational expression for his average speed.
B. On Tuesday he drove for 6 hours at the same average speed. Write a rational expression for his distance on Tuesday.
I haven't a clue how to set this up or do it.Can someone help me?Thanks in advance.

Answer by stanbon(75887)

Henry drove 800 miles in X hours on Monday.
A. Write a rational expression for his average speed.
average speed = distance/time = (800/x) miles per hour
--------------------------------------------------------------
B. On Tuesday he drove for 6 hours at the same average speed. Write a rational expression for his distance on Tuesday.
distance = rate * time = [(800/x)*6] = (4800/x) miles
=====================================
Cheers,
Stan H.